OpenStudy (anonymous):

What two numbers added equals -5 and multiplied equals positive 3

OpenStudy (mathmath333):

\[a+b=-5\]\[a*b=3\]

OpenStudy (mathmath333):

u can substitute \[a=3/b\] in first equation

OpenStudy (anonymous):

so how do I start?

OpenStudy (anonymous):

-.6

OpenStudy (mathmath333):

\[3/b+b=-5\]

OpenStudy (mathmath333):

solve it for

OpenStudy (mathmath333):

b

OpenStudy (anonymous):

b+b= -15?

OpenStudy (anonymous):

Multiplied 3 to both sides

OpenStudy (mathmath333):



OpenStudy (mathmath333):

multilply it by b

OpenStudy (mathmath333):

u know BODMAS rule

OpenStudy (anonymous):

no but I know pemdas?

OpenStudy (mathmath333):

ya its the same

OpenStudy (mathmath333):

it says division comes first then addition

OpenStudy (anonymous):

-2

OpenStudy (mathmath333):

so u can't add b+b and multilply 5and3

OpenStudy (anonymous):

1b

OpenStudy (mathmath333):

first u have to solve the division

OpenStudy (anonymous):

-5/3

OpenStudy (mathmath333):